OpenStudy (anonymous):

The table below shows two equations: Equation 1 |2x – 3| + 5 = 4 Equation 2 |5x + 3| − 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 2, -3.2. The solutions to equation 1 are x = 1, 2 and equation 2 has no solution. The solutions to equation 1 are x = 1, 2 and equation 2 has solutions x = 2, -3.2.

OpenStudy (anonymous):

mode>=0 always first |2x+3|=4-5=-1 not possible so no solution

OpenStudy (anonymous):

second |5x+3|=13 5x+3=pm13

OpenStudy (anonymous):

5x+3=+-13 5x+3=13 and 5x+3=-13 5x=10 5x=-16 x=2 x=-16/5

OpenStudy (anonymous):

its the second answer?

OpenStudy (anonymous):

It's c :)